# arXiv:2603.00774v1[cs.HC] 28 Feb 2026

### v1

@@ Abstract.
$c$s: Human-centered computing Empirical studies in HCIccs: Human-centered computing Natural language interfacesccs: Applied computing Health care information systems
@@ 1. Introduction
Conversational agents are increasingly explored as scalable mental health support tools, offering on-demand access, anonymity, and reduced cost compared to traditional care (Jabir et al., 2023; Kang and Hong, 2025; Tong et al., 2025; He et al., 2023a). Advances in large language models (LLMs), such as ChatGPT, now enable open-ended, context-sensitive conversations, expanding the expressive capacity of mental health chatbots (Lee et al., 2020; Moell, 2024). At the same time, structured therapeutic frameworks can be embedded into conversational systems. The Self-Attachment Technique (SAT) (Edalat, 2015) is a self-administered, attachment-based approach that supports emotion regulation and self-compassion by cultivating an imaginative affectional bond between the adult self and the childhood self of an individual (Alazraki et al., 2021; Elahimanesh et al., 2023). Prior SAT chatbots using rule-based or classification-driven dialogue demonstrated promising empathy and engagement, but were limited in flexibility and scalability (Alazraki et al., 2021; Elahimanesh et al., 2023).
@@
Despite these advances, most mental health chatbots emphasize feasibility and acceptability while relying on shallow interaction patterns and short-term evaluations, leaving open questions about how system architecture shapes perceived therapeutic quality (Kang and Hong, 2025; Tong et al., 2025; He et al., 2023a). Research highlights the importance of personalization, empathic responses, and sustained engagement (He et al., 2023a; Jabir et al., 2023), yet existing work rarely compares architectural paradigms under controlled conditions (Zhang et al., 2025). Recent proposals such as multi-agent orchestration and finite state machines offer ways to structure complex LLM behavior (Zhang et al., 2025), but their impact on therapeutic naturalness, authenticity, and engagement, particularly for structured interventions like SAT, remains underexplored.
@@
To address these gaps, we conducted a between-subjects randomized controlled trial (RCT) examining how LLM-based architectural design influences perceived therapeutic quality in a SAT-informed chatbot. Using GPT-4o (OpenAI et al., 2024) as the base model across all conditions, we compared three architectures. The Intervention employed a multi-agent finite state machine aligned with SAT stages, incorporating explicit state transitions, a SAT knowledge base, and shared long-term memory among the agents. The Active Control used a single agent with equivalent prompts and SAT access. The Placebo Control used GPT-4o without SAT knowledge or structure, resembling a generic conversational companion. System architecture and prompts were authored in English, while deployment targeted a Farsi-speaking user population.
@@
We deployed the system for eight days with $N=66$ participants (Intervention: 22, Control: 23, Placebo: 21), who completed demographic measures, interacted with their assigned chatbot, and completed post-study surveys. The intervention was perceived as significantly more natural and human-like than both control conditions (F=7.017, p=0.0018, $\eta^{2}$=0.187), with mostly higher but non-significant scores on other therapeutic metrics, such as trust, memory coherence, user satisfaction, fewer off-topic responses, and ease of use. Chat logs showed the intervention produced more but shorter messages (459 messages, $\sim$230 characters) than the active control (336 messages, $\sim$409 characters) and placebo control (206 messages, $\sim$635 characters), indicating a more interactive dialogue style. Findings demonstrate that architectural choices, beyond prompt design alone, meaningfully shape user experience in LLM-based therapeutic chatbots.
@@ 3. Method
System Design and Architecture. To investigate how conversational structure and adaptivity influence therapeutic engagement, we developed a three-tier conversational AI system that delivers the Self-Attachment Technique within a Randomized Controlled Trial framework. The study included three experimental conditions: (1) a finite state machine (FSM) driven multi-agent system with long-term memory intervention (Alpha), (2) a single-agent system with a knowledge-base as an active control (Beta), and (3) a minimal single-agent placebo (Gamma). All versions run on a Django REST backend in Python and employ OpenAI’s GPT-4o (OpenAI et al., 2024) as the conversational engine. User interactions occurred in Farsi, supported by a specialized response retrieval system for intent classification, while prompts and conversational structures were authored in English. To maximize ecological validity, participants accessed the intervention via a React.js web interface, identical across all three conditions.
@@
Experimental Conditions and Randomization. Upon registration, participants were randomly assigned to one of three groups and blinded to their condition to minimize expectancy effects. All systems employed the same LLM and shared interface, response style, and interaction modality. (1) Alpha implements a 12-state therapeutic protocol based on the SAT stages (Elahimanesh et al., 2023), while supporting open-ended conversation throughout the flow. At each stage, dedicated agents guide state transitions using LLM-driven decision logic, while maintaining open-ended dialogue to engage users and smoothly navigate the conversation to the next stage. To ensure consistency among the agents, the system utilizes a shared long-term conversational memory, generated via prompt-guided LLM summarization to preserve emotional states, personal context, and salient experiences. Additionally, it incorporates an agent using an adaptive Retrieval-Augmented Generation (RAG) mechanism to deliver personalized therapeutic exercises; (2) Beta delivers the identical SAT content and daily exercises but without FSM enforcement, relying solely on a single comprehensive system prompt to implicitly guide the conversation through similar phases within the LLM’s context window. To isolate the effect of architecture, this version uses the exact same prompts as Alpha, collapsed into a single sequential prompt; (3) Gamma provides a minimal conversational single-agent, controlling for the “digital companion” effect. It uses a stripped-down prompt with no SAT knowledge, no exercise recommendation engine, and no structured therapeutic goals, representing a near–pure instantiation of the LLM.
@@
Conversational Structure: The Finite State Machine. A key contribution of the Alpha system is its explicit modeling of the therapeutic process. The FSM divides the interaction into four logical Superstates or phases: Initiation (Greeting/Emotion Assessment), Exploration (Event Processing), Intervention (Exercise Delivery), and Conclusion. Unlike rigid rule-based chatbots that rely on keyword matching, the Alpha system employs two advanced transition mechanisms to maintain naturalness: (1) LLM-as-Judge Sufficiency Detection (Zheng et al., 2023a; Szymanski et al., 2025), where a dedicated prompt evaluates whether accumulated user responses satisfy emotional and informational criteria for the current state, ensuring meaningful expression of the feelings and events before state change; (2) Intent-Based Routing, which classifies over 40 affirmative/negative Farsi linguistic patterns to route the user to the correct next state.
@@ 4. Experiment and Results
The Primacy of Conversational Naturalness. The strongest finding was a significant architectural effect on perceived naturalness and human-like tone. Alpha was rated higher ($\mu=3.955$, $SD=0.950$) than Beta ($\mu=3.043$, $SD=0.825$) and Gamma ($\mu=3.211$, $SD=0.787$) ($F=7.017$, $p_{perm}=0.0018$, $\eta^{2}=0.187$), reflecting a nearly one-point gain on a 5-point scale (0.912 over Beta) and a shift from “somewhat” to “quite” natural. The multi-agent FSM architecture explained $\sim$19% of rating variance, a moderate-to-large effect. Pairwise tests showed Alpha outperformed both Beta and Gamma, which were similar, indicating benefits of explicit multi-agent state management beyond content alone. All conditions showed comparable baseline usability (ease of use: $F=0.204$, $p=0.826$, $\eta^{2}=0.007$; feature integration: $F=0.859$, $p=0.446$, $\eta^{2}=0.027$), isolating improvements to relational quality rather than functional UX.
Table 1. Survey results compared three architectures on seven interaction metrics (1–5 Likert). Alpha almost consistently outperformed Beta and Gamma, with higher mean values. For off-topic replies, lower scores are better. Naturalness was statistically significant. F-statistics show variance between versus within groups, and $\eta^{2}$ indicates the proportion of variance explained by architecture.
@@
Evaluated Interaction Metrics. As shown in Table 1, Alpha trended higher in the majority of other experiential metrics as well, including memory for past conversations ($\mu=3.682$ vs. 3.348/3.263), perceived empathy ($\mu=3.727$ vs. 3.435/3.895), and trust, while these differences were not statistically significant. Alpha also reported fewer off-topic replies ($\mu=1.727$ vs. 2.304/2.000), indicating better conversational focus, though not significant ($F=1.481$, $p=0.225$, $\eta^{2}=0.046$). Overall satisfaction was moderately positive across groups ($\mu\approx 3.4$–3.6, $F=0.520$, $p=0.606$, $\eta^{2}=0.017$), showing general acceptability regardless of architecture. These patterns suggest that architectural design mainly drives perceived naturalness and conversational dynamics, while also enhancing relational qualities like trust and empathy; detecting significance for these secondary effects may require larger samples or distinct interventions.
@@
In addition, importantly the FSM architecture in Alpha, explicitly structured around therapeutic stages, naturally reflects the reasoning of a human therapist, guiding the conversation through problem identification, exploration, and resolution. Multi-agent coordination keeps each agent focused on its stage, while long-term memory preserves context within and across sessions, enhancing continuity and relevance. A structured therapeutic knowledge base ensures contextually appropriate guidance. Beta’s prompting and knowledge contributed to its intermediate performance, and Gamma’s minimal design performed lowest. Together, these elements suggest that Alpha’s design, by mimicking human conversational dynamics, multi-agent specialization, memory retention, and grounded expertise, directly drives higher perceived naturalness and overall interaction quality.
@@ 5. Discussion
Our findings show that architectural choices in LLM-based therapeutic chatbots strongly shape user experience, beyond prompt engineering and toward fundamental questions of system structure, due to the staged, process-driven nature of therapy. Multi-agent architectures with explicit state management produced more natural interactions than single-agent or unguided LLMs, indicating that effective chatbots require careful orchestration alongside content delivery.
@@
Conversational Naturalness. The intervention’s advantage in perceived naturalness reflects more than stylistic preference. Natural dialogue aligns with person-centered care, where interaction quality predicts outcomes (Miller and Rollnick, 2012). Chat logs show that the multi-agent system produced shorter, more interactive exchanges, shifting from monologic responses to genuine dialogue and challenging assumptions that chatbots should maximize information per turn. Instead, effective conversational AI should mirror human dialogue rhythms through frequent turn-taking and collaborative meaning-making. By distributing therapeutic stages across specialized agents, the architecture likely enabled this dialogic pattern, with long-term memory further contributing by preserving users’ feelings and experiences across sessions. Notably, gains in naturalness did not extend to empathy, trust, or personalization, suggesting that these are related yet distinct dimensions of the therapeutic alliance. Architectural improvements that enhance naturalness alone may therefore be insufficient to foster emotional attunement, which may depend more strongly on prompt design or response content.
@@
Structural Scaffolding and Design Implications. Across most interaction metrics, including naturalness, trust, memory coherence, user satisfaction, off-topic responses, and ease of use, performance consistently followed Alpha >> Beta >> Gamma. This pattern highlights the role of therapeutic structure and architectural scaffolding in shaping perceived interaction quality, even without user awareness of system design. Stronger structural guidance supports coherent dialogue, relevant questioning, and alignment with therapeutic goals, without limiting conversational flexibility.
@@
Limitations and Future Directions. Our study highlights several limitations and directions for future work: (1) the eight-day deployment captured only early user impressions, leaving long-term engagement, outcomes, and the development of user-agent relationships unresolved, including whether perceived naturalness translates into sustained use or longer-term benefits; (2) the use of single-item measures for empathy, trust, and personalization limits construct validity, and future studies should employ validated multi-item instruments (e.g., the Working Alliance Inventory (Hatcher and Gillaspy, 2006)) and examine how architectural effects vary across user characteristics; (3) the focus on Persian-speaking, highly educated participants (93.9% holding at least a Bachelor’s degree) and an English-authored system architecture constrains cross-linguistic and broader population generalizability; (4) therapeutic outcomes and skill acquisition were not directly assessed, underscoring the need to link subjective interaction quality to objective behavioral and clinical measures; and (5) the collected chatbot interaction data represent a valuable therapeutic dataset that, when combined with other sources, could support LLM alignment using methods such as PPO or DPO (Zheng et al., 2023b; Rafailov et al., 2024).

### v2

# Structure Matters: Evaluating Multi-Agents Orchestration in Generative Therapeutic Chatbots

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Natural language interfacesccs: Applied computing Health care information systems
@@ 1. Introduction
Conversational agents are increasingly explored as interactive systems for supporting sensitive domains such as mental health, offering on-demand access, anonymity, and reduced cost compared to traditional care (Jabir et al., 2023; Kang and Hong, 2025; Tong et al., 2025; He et al., 2023a). Advances in large language models (LLMs), such as ChatGPT, now enable open-ended, context-sensitive conversations, expanding the expressive capacity of mental health chatbots (Lee et al., 2020; Moell, 2024). At the same time, structured therapeutic frameworks can be embedded into conversational systems. The Self-Attachment Technique (SAT) (Edalat, 2015) is a self-administered, attachment-based approach that supports emotion regulation and self-compassion by cultivating an imaginative affectional bond between the adult self and the childhood self of an individual (Alazraki et al., 2021; Elahimanesh et al., 2023). Prior SAT chatbots using rule-based or classification-driven dialogue demonstrated promising empathy and engagement, but were limited in flexibility and scalability (Alazraki et al., 2021; Elahimanesh et al., 2023).
@@
Despite these advances, most mental health chatbots emphasize feasibility and acceptability, with comparatively limited attention to the underlying conversational design structure and interaction architecture, leaving open questions about how different architectural paradigms shape user experience  (Kang and Hong, 2025; Tong et al., 2025; He et al., 2023a). Research highlights the importance of personalization, empathic responses, and sustained engagement (He et al., 2023a; Jabir et al., 2023), yet existing work rarely compares architectural paradigms under controlled conditions (Zhang et al., 2025). Recent proposals such as multi-agent orchestration and finite state machines offer ways to structure complex LLM behavior (Zhang et al., 2025), but their impact on perceived conversational qualities such as naturalness, trust and engagement, particularly for structured interventions remains underexplored.
@@
To address these gaps, we conducted a between-subjects randomized controlled trial (RCT) examining how LLM-based architectural design influences users’ perceived conversational and interaction quality in a SAT-informed chatbot. Using GPT-4o (OpenAI et al., 2024) as the base model across all conditions, we compared three architectures. The Intervention employed a multi-agent finite state machine inspired by SAT stages, incorporating explicit state transitions, a SAT knowledge base, and shared long-term memory among the agents. The Active Control used a single agent with equivalent prompts and SAT knowledge base. The Placebo Control used GPT-4o without SAT knowledge or structure, resembling a generic conversational companion. System architecture and prompts were authored in English, while deployment targeted a Farsi-speaking user population.
@@
We deployed the system for eight days with $N=66$ participants (Intervention: 22, Control: 23, Placebo: 21), who completed demographic measures, interacted with their assigned chatbot, and completed post-study surveys. The intervention was perceived as significantly more natural and human-like than both control conditions, with mostly higher but non-significant scores on other self-reported user experience measures, such as trust, memory coherence, user satisfaction, fewer off-topic responses, and ease of use. Chat logs showed the intervention produced more but shorter messages (459 messages, $\sim$230 characters) than the active control (336 messages, $\sim$409 characters) and placebo control (206 messages, $\sim$635 characters), indicating differences in turn-taking patterns and conversational pacing. Overall, this work contributes an empirical comparison of architectural design patterns for LLM-based conversational user interfaces (CUIs), showing how system-level structure shapes conversational interaction and user experience. The evaluation is based on interaction-level metrics from a non-clinical population and does not address clinical outcomes or therapeutic efficacy. These findings suggest implications for how architectural design and memory mechanisms can affect conversational systems for sensitive domains such as mental health.
@@ 3. Proposed Method
System Design and Architecture. To investigate how conversational structure and adaptivity influence therapeutic engagement, we developed a three-tier conversational AI system that delivers the Self-Attachment Technique within a Randomized Controlled Trial framework. The study included three experimental conditions: (1) a finite state machine (FSM) driven multi-agent system with long-term memory intervention (Alpha), (2) a single-agent system with a knowledge-base as an active control (Beta), and (3) a minimal single-agent placebo (Gamma). All versions run on a Django REST backend in Python and employ OpenAI’s GPT-4o (OpenAI et al., 2024) as the conversational engine. User interactions occurred in Farsi, supported by a specialized response retrieval system for intent classification, while prompts and conversational structures were authored in English. To maximize ecological validity, participants accessed the intervention via a React.js web interface, identical across all three conditions. (see Appendix B for a screenshot).
@@
Experimental Conditions and Randomization. Upon registration, participants were randomly assigned to one of three groups and blinded to their condition to minimize expectancy effects. All systems employed the same LLM and shared interface, response style, and interaction modality. (1) Alpha implements a 12-state therapeutic protocol based on the SAT stages (Elahimanesh et al., 2023), while supporting open-ended conversation throughout the flow. (see Appendix A for stage details). At each stage, dedicated agents guide state transitions using LLM-driven decision logic, while maintaining open-ended dialogue to engage users and smoothly navigate the conversation to the next stage. To ensure consistency among the agents, the system utilizes a shared long-term conversational memory, generated via prompt-guided LLM summarization to preserve emotional states, personal context, and salient experiences. Additionally, it incorporates an agent using an adaptive Retrieval-Augmented Generation (RAG) mechanism to deliver personalized therapeutic exercises; (2) Beta delivers the identical SAT content and daily exercises but without FSM enforcement, relying solely on a single comprehensive system prompt to implicitly guide the conversation through similar phases within the LLM’s context window. To isolate the effect of architecture, this version uses the exact same prompts as Alpha, collapsed into a single sequential prompt; (3) Gamma provides a minimal conversational single-agent, controlling for the “digital companion” effect. It uses a stripped-down prompt with no SAT knowledge, no exercise recommendation engine, and no structured therapeutic goals, representing a near–pure instantiation of the LLM.
@@
Conversational Structure: The Finite State Machine. A key contribution of the Alpha system is its explicit modeling of the therapeutic process. It is operated using a programmatic FSM, based on the SAT flowchart (Elahimanesh et al., 2023). The FSM encompasses 12 distinct states and divides the interaction into four logical Superstates or phases: Initiation (Greeting/Emotion Assessment), Exploration (Event Processing), Intervention (Exercise Delivery and Feedback Loop), and Conclusion. Unlike rigid rule-based chatbots that rely on keyword matching, the Alpha system employs two advanced transition mechanisms to maintain naturalness: (1) LLM-as-Judge Sufficiency Detection (Zheng et al., 2023a; Szymanski et al., 2025), where a dedicated prompt evaluates whether accumulated user responses satisfy emotional and informational criteria for the current state, ensuring meaningful expression of the feelings and events before state change; (2) Intent-Based Routing, which classifies over 40 affirmative/negative Farsi linguistic patterns to route the user to the correct next state.
@@ 4. Experiment and Results
Conversational Naturalness as Primary Effect. The strongest finding was a significant architectural effect on perceived naturalness and human-like tone. Alpha was rated higher ($\mu=3.955$, $SD=0.950$) than Beta ($\mu=3.043$, $SD=0.825$) and Gamma ($\mu=3.211$, $SD=0.787$) ($F=7.017$, $p_{perm}=0.0018$, $\eta^{2}=0.187$), reflecting a nearly one-point gain on a 5-point scale (0.912 over Beta) and a shift from “somewhat” to “quite” natural. This corresponds to $\sim$19% of variance in perceived naturalness, indicating a moderate-to-large effect associated with the Alpha system design. Pairwise tests showed Alpha outperformed both Beta and Gamma, which were similar, indicating that the structured Alpha design was associated with more natural conversational experiences than the comparison conditions. All conditions showed comparable baseline usability (ease of use: $F=0.204$, $p=0.826$, $\eta^{2}=0.007$; feature integration: $F=0.859$, $p=0.446$, $\eta^{2}=0.027$), suggesting that observed differences primarily concern perceived conversational quality rather than functional usability.
Table 1. Survey results compared three architectures on seven interaction metrics (1–5 Likert). Alpha generally achieved higher mean values than Beta and Gamma across several metrics. For off-topic replies, lower scores are better. Naturalness was statistically significant. F-statistics show variance between versus within groups, and $\eta^{2}$ indicates the proportion of variance explained by architecture.
@@
Evaluated Interaction Metrics. In this study, we evaluated interaction-level user experience and conversational qualities using self-reported metrics. As shown in Table 1, Alpha trended higher in the majority of other experiential metrics as well, including memory for past conversations ($\mu=3.682$ vs. 3.348/3.263), perceived empathy ($\mu=3.727$ vs. 3.435/3.895), and trust, while these differences were not statistically significant. Alpha also reported fewer off-topic replies ($\mu=1.727$ vs. 2.304/2.000), indicating better conversational focus, though not significant ($F=1.481$, $p=0.225$, $\eta^{2}=0.046$). Overall satisfaction was moderately positive across groups ($\mu\approx 3.4$–3.6, $F=0.520$, $p=0.606$, $\eta^{2}=0.017$), showing general acceptability regardless of architecture. These patterns suggest that architectural design mainly influences perceived naturalness and conversational dynamics, while evidence for effects on relational qualities such as trust and empathy remains inconclusive under the present sample size.
@@
In addition, importantly the FSM architecture in Alpha, explicitly structured around therapeutic stages, naturally reflects the reasoning of a human therapist, guiding the conversation through problem identification, exploration, and resolution. Multi-agent coordination keeps each agent focused on its stage, while long-term memory preserves context within and across sessions, enhancing continuity and relevance. A structured therapeutic knowledge base ensures contextually appropriate guidance. Beta’s prompting and knowledge contributed to its intermediate performance, and Gamma’s minimal design performed lowest. Together, these elements suggest that Alpha’s design, by mimicking human conversational dynamics, multi-agent specialization, memory retention, and grounded expertise, is associated with higher perceived naturalness and improved conversational dynamics.
@@ 5. Discussion
Our findings show that architectural choices in LLM-based conversational agents strongly shape perceived interaction quality, beyond prompt design and toward fundamental questions of system structure, due to the staged and process-driven nature of therapy. Multi-agent architectures with explicit state management were associated with more natural interactions than single-agent or unguided LLMs, suggesting that effective CUI design require careful orchestration alongside content delivery.
@@
Conversational Naturalness. Observed advantages in perceived naturalness reflect more than stylistic preference. Chat logs indicate that the multi-agent system produced shorter, more interactive exchanges, shifting from monologic responses to dialogue-like interactions and challenging the assumption that chatbots should maximize information per turn. Effective CUI design may therefore benefit from mirroring human dialogue rhythms, such as frequent turn-taking and collaborative meaning-making. By distributing therapeutic stages across specialized agents, the architecture likely enabled this dialogic pattern, with long-term memory further contributing by preserving users’ feelings and experiences across sessions. Notably, gains in naturalness did not extend to empathy, trust, or personalization, suggesting that these are related yet distinct dimensions of the therapeutic alliance. Architectural improvements that enhance naturalness alone may therefore be insufficient to foster emotional attunement, which may depend more strongly on prompt design or response content.
Structural Scaffolding and Design Implications.
@@
Across several interaction measures, Alpha generally outperformed one or both comparison conditions, with the clearest and only statistically reliable effect observed for naturalness. This pattern highlights the role of therapeutic structure and architectural scaffolding in shaping perceived interaction quality in CUIs, even without user awareness of system design. Stronger structural guidance supports coherent dialogue, relevant questioning, and alignment with therapeutic goals, without limiting conversational flexibility. Additionally, these implications can inform the design of conversational agents in other sensitive domains, such medical contexts.
@@
Limitations and Future Directions. Our study has several limitations and suggests directions for future work: (1) The eight-day deployment captured only early user impressions, leaving long-term engagement, outcomes, and the development of user-agent relationships unresolved, including whether perceived naturalness translates into sustained use or longer-term benefits. (2) The use of single-item measures for empathy, trust, and personalization limits construct validity; future studies should employ validated multi-item instruments (e.g., the Working Alliance Inventory (Hatcher and Gillaspy, 2006)) and examine how architectural effects vary across user characteristics. (3) The focus on Persian-speaking, educated participants (93.9% holding at least a Bachelor’s degree) constrains cross-linguistic and broader population generalizability. Future work could explore more diverse linguistic, cultural, and educational populations. (4) No clinical or therapeutic outcomes were assessed, and findings reflect perceived interaction quality rather than efficacy. Future work can investigate the effect of interaction metrics on therapy efficiency in a clinical population. (5) The bundled design (multi-agent orchestration and memory) prevents attributing effects to specific components. Because in the multi-agent variation there must be mechanisms for information sharing among agents, it was not possible to implement it without a shared memory. Thus, in the intervention version, multi-agent architecture and memory coexist, which represents a limitation of our work. While we preserved exactly equal prompts and knowledge bases for the alpha and beta versions, the findings indicate that multi-agent architecture combined with memory and knowledge base performed better compared to a single-agent system with the same knowledge base, and also outperformed a general unguided LLM agent. Future work can investigate each of these components separately through an ablation study. (6) The collected chatbot interaction data represent a valuable therapeutic dataset that, when combined with other sources, could support LLM alignment using methods such as PPO or DPO (Zheng et al., 2023b; Rafailov et al., 2024).